### v2

# HD-GEN: A High-Performance Software System for Human Mobility Data Generation Based on Patterns of Life

Abstract.
Understanding individual-level human mobility is critical for a wide range of applications. As such, real-world trajectory datasets provide valuable insights into actual movement behaviors and patterns of life but are often constrained by data sparsity and participant bias. Synthetic data, by contrast, offers scalability and flexibility but frequently lacks realism. To address this gap, we introduce a comprehensive software pipeline for, generating, calibrating, processing, and visualizing large-scale individual-level human mobility datasets that combine the realism of empirical data with the control and extensibility of Patterns-of-Life simulations. Our system consists of four integrated components. (1) a data generation engine which constructs geographically grounded simulations using OpenStreetMap data to produce diverse mobility logs. (2) a genetic algorithm–based calibration module that fine-tunes simulation parameters to align with real-world mobility characteristics, such as daily trip counts and radius of gyration, enabling realistic behavioral modeling. (3) a data processing suite which transforms raw simulation logs into structured formats suitable for downstream applications, including model training and benchmarking, and (4) a visualization module that extracts key mobility patterns and insights from the processed datasets and presents them through intuitive visual analytics for improved interpretability.
Trajectory Generation, Patterns of Life, Simulation, Trajectory Data
1. Introduction
Human mobility datasets are fundamental for understanding human behavior (Toch et al., 2019; Zhu et al., 2024a; Gonzalez et al., 2008), mobility patterns (Mokbel et al., 2024; Schneider et al., 2013), and traffic dynamics (Chen et al., 2022; Islam et al., 2021). They enable a wide range of applications, including outlier and anomaly detection (Zhang et al., 2024a, b; Amiri et al., 2024b), infectious disease modeling (Kohn et al., 2023), and urban planning and infrastructure design (Isaacman et al., 2012). Mobility datasets can be broadly categorized as real or synthetic. Real world data, such as those collected from cell phone traces, provide high fidelity but raise serious quality and privacy concerns, while synthetic data offer improved scalability and privacy preservation at the cost of reduced realism (Züfle et al., 2024). In practice, privacy regulations (Xiao and Xiong, 2015; Khoshgozaran et al., 2011; Krumm, 2009), de-identification techniques, and voluntary data collection introduce sparsity and information loss, resulting in significant trade-offs between data utility, realism, and accessibility (Zhu et al., 2024b; Amichi et al., 2025).
To mitigate these limitations, researchers are increasingly relying on simulated trajectory datasets which provide scalable, noise free mobility information without privacy risks. However, many of the existing trajectory simulators rely on random or simplistic destination selection (Düntgen et al., 2009; Brinkhoff, 2002) or parametric trip distributions (Levandoski et al., 2012; Armenatzoglou et al., 2013). While such approaches are suitable for benchmarking storage systems or query processing, we would argue that they fail to capture the intentional and routine driven nature of real human mobility.
The Pattern of Life (POL) simulation framework (Kim et al., 2020; Züfle et al., 2023; Amiri et al., 2024a) addresses this gap by modeling human needs, activities, and daily routines. POL generates purposeful mobility behaviors such as commuting, dining, and social interactions, producing trajectories that better reflect real world movement semantics (e.g., going to and from work). POL has been used to generate large scale datasets with realistic mobility patterns, social interactions, and check-in behavior. Prior work includes the generation of fine grained mobility and social network datasets across multiple U.S. cities (Amiri et al., 2023), datasets with injected anomalies for mobility anomaly detection (Amiri et al., 2024b), and a GeoLife calibrated dataset optimized via genetic algorithms to balance realism and scalability (Amiri et al., 2024c).
In this work, we extend our SIGSPATIAL 2025 short paper (Amiri et al., 2025) and consolidate our prior contributions (Amiri et al., 2024a, c, 2023) into a complete and systematic POL based data generation system. We introduce a unified software platform that integrates trajectory generation, calibration, processing, and visualization into a coherent end to end pipeline, referred to as HD-GEN (i,e., Human mobility Data GENeration based on POL) . The platform expands the functionality of POL while emphasizing reproducibility, scalability, and usability for large scale experimental and operational settings.
Building on the short paper, this extended version makes the following additional contributions:
  • improves reproducibility by introducing checkpointing mechanisms, providing practical guidance on their use, and supplying scripts for efficient checkpoint management
  • provides a detailed description of the logging infrastructure and introduces a streaming framework to support real time analytics and monitoring workflows
  • enhances scalability for large scale parallel execution, enabling efficient simulation on distributed and multi core systems, and supporting calibration and checkpoint based execution strategies
  • improves usability through Python based tools that simplify configuration and execution, reducing reliance on low level command line interfaces and manual scripting
  • increases interpretability by adding a visualization phase that produces actionable insights into simulated mobility patterns and system behavior
  • delivers an improved open source repository ({https://github.com/onspatial/hd-gen}) with clearer documentation and integrated visualization toolkits.
The remainder of the paper is organized as follows. In Section 2, we review relevant background information and related prior work. Next , in Section 3, we provide a technical overview of POL and describe the software design and system-level details of HD-GEN. In Section 4, we summarize the datasets produced using HD-GEN. Finally, in Section 5, we conclude by summarizing the contributions of this work and discussing directions for future research.
2. Background Information and Related Work
In this section, we provide background information and related work on human mobility data generation, including an overview of the POL simulation and a discussion of existing trajectory datasets. We first describe the POL framework, its behavioral foundations, and its suitability for generating large-scale, realistic mobility data. We then briefly review commonly used real and synthetic trajectory datasets and highlight their limitations, motivating the use of POL and the need for calibration against observed mobility patterns.
2.1. The Patterns-of-Life Simulation (POL)
The POL simulation, introduced in Urban Life (Züfle et al., 2023), is a city-scale, agent-based modeling framework designed to generate realistic human mobility trajectories and interaction patterns. The simulation instantiates autonomous agents within an urban environment constructed from OpenStreetMap data, including road networks, residential zones, workplaces, and recreational venues. Agents perceive and interact with this environment and each other while independently making decisions that reflect the routines and choices of individuals in everyday life.
Agent behavior in POL is governed by a combination of physiological, financial, and social needs inspired by Maslow’s hierarchy of needs (Maslow, 1943), together with goal-oriented decision making based on the Theory of Planned Behavior (Ajzen, 1991). At each simulation time step, agents evaluate competing needs such as earning income, acquiring food, maintaining social relationships, and participating in leisure activities. These needs are balanced against practical constraints including time availability, spatial accessibility, and personal resources. The resulting decisions determine when and where agents travel, producing realistic daily schedules and mobility patterns. As agents move between different types of locations, the simulation generates fine-grained spatiotemporal trajectories as well as implicit social graphs that emerge from repeated co-location and interaction. This design makes POL well suited for controlled experimentation and downstream analysis in mobility modeling, social network analysis, and urban studies.
Figure 1. Interaction between a person agent and the simulation environment for generating human mobility logs.
Fig. 1 illustrates the interaction between a person agent and the simulation world. Each person is modeled as an autonomous agent with multiple needs, including food, sleep, shelter, financial safety, and love need from social connections. These needs evolve continuously over time and drive the agent’s decision making process. To satisfy its needs, the agent interacts with the environment, which defines mobility constraints, a spatial network, and points of interest such as workplaces, homes, and public venues. The environment determines where the agent can move and which actions are feasible at each location. Each agent perceives their environment, selects actions based on its current needs and context, and executes those actions through movement and interactions with points of interest, such as working at a workplace, or with other agents at recreational sites. All agent states, actions, and environment responses are recorded by the log system. The generated logs capture realistic spatiotemporal behavior and serve as the output of the simulation for downstream analysis and modeling.
Figure 2. End to end POL simulation pipeline showing parameter configuration, map generation, initialization, execution, checkpointing, and structured log generation
The POL simulation is implemented in Java utilizing the MASON toolkit (Luke et al., 2018) and is configurable through a command-line interface, and publicly available on GitHub (Züfle et al., 2023). Users can specify city regions, population sizes, and behavioral parameters, enabling reproducible and extensible experiments. Fig. 2 presents the end to end simulation workflow. The process begins with a parameter sweep, where multiple configurations are defined to explore variations in agent behavior and environmental conditions. These configurations are formalized during parameterization, producing a consistent set of parameters that govern the simulation. In parallel, the map generator produces spatial representations of the environment, including networks and points of interest. The generated maps, together with the selected parameters, are passed to the initialization stage, where agents, environments, and initial states are instantiated. The simulation stage executes agent interactions over time based on the initialized state and active parameters. During execution, manipulation modules can dynamically modify parameters or system states, enabling interventions, scenario changes, or policy testing. The simulation state can be serialized to disk to implement checkpointing, allowing execution to be paused, restored, and reproduced in long running experiments. All simulation outputs are stored in structured formats guided by a predefined schema definition, which specifies how agent states, actions, and environmental events are recorded. The log system collects these records, producing standardized logs that serve as the final output of the simulation.
In this system paper, we build directly on the original POL framework and preserve its behavioral and decision making logic while introducing new features and targeted improvements in execution efficiency and scalability. These optimizations enable simulations with larger populations and longer time horizons without altering the underlying semantics of agent behavior, ensuring comparability with prior work while substantially improving practical usability.
2.2. Related Work and Relation to HD-GEN
Human mobility data can be collected from real-world observations or synthesized using computational models. Real-world trajectory datasets capture observed mobility behavior and provide valuable insight into human movement and activity patterns. A widely used example is the GeoLife dataset (Zheng et al., 2011), which contains four years of high-resolution GPS traces from approximately 180 users in Beijing, spanning both commuting and leisure activities. More recently, the large-scale YJMOB100K dataset (Yabe et al., 2024) provides anonymized mobility records from 100,000 users at 30-minute intervals, aggregated into 500-meter spatial cells. While this scale enables population-level analysis, the spatial aggregation limits fine-grained, place-based behavioral inference. Additional trajectory datasets focus on vehicles, including taxis (Yuan et al., 2010; Piorkowski et al., 2009) and buses (Dias and Costa, 2018). These datasets are valuable for modeling transportation systems but do not accurately represent individual human activity patterns, since each trip may involve different occupants and does not reflect continuous personal behavior. Synthetic trajectory datasets address limitations of real mobility data, including privacy constraints, proprietary access, and restricted scale. Recent deep generative approaches learn mobility patterns to produce realistic spatiotemporal trajectories, including factorized models such as EETG (Zhang et al., 2022) and diffusion-based methods (Zhu et al., 2024b). Optimization-based techniques further aim to enhance incomplete or sparse trajectory data (Zhu et al., 2024a; Hu et al., 2022). In contrast, the POL simulation framework (Kim et al., 2020; Züfle et al., 2023; Amiri et al., 2024a) generates trajectories driven by modeled human needs and decision processes rather than purely geometric or kinematic realism.
In our prior work (Amiri et al., 2023), POL was used to generate a large-scale dataset exceeding 1.5 TB of mobility, check-in, and social interaction data. We further demonstrated the resulting data generator and its ability to reproduce key mobility and activity patterns in (Amiri et al., 2024a). Although the default parameter settings yield qualitatively realistic behavior, regional variation was not systematically characterized. This limitation motivated our subsequent study (Amiri et al., 2024c), in which we calibrated POL to better match mobility patterns observed in the GeoLife dataset for Beijing. Finally, we integrated these efforts into a unified framework, HD-GEN, which was introduced in a concise system paper (Amiri et al., 2025).
3. Software Architecture and System Details
This section describes the overall software architecture and execution flow of the proposed system. We first outline the software architecture, highlighting the main components and their interactions. We then detail the POL simulation, focusing on its technical design, scenario based testing with checkpointing, and real time log streaming. Finally, we present the HD-GEN pipeline, describing each phase of the workflow from data generation to calibration, processing, and visualization, and explaining how these stages integrate into a coherent end to end pipeline.
3.1. Software Architecture
Our system enables the generation of synthetic datasets that either statistically replicate real-world mobility patterns or conform to user-defined specifications. Fig. 3 illustrates the architecture of the proposed software. The dataset generation pipeline is organized into four main phases: Generation, Calibration, Processing, and Visualization. The figure highlights the relationships among these phases and their interaction within the overall software framework. Each phase has its own internal sub-architecture, which is described in detail in Section 3.4.
Figure 3. Software architecture of HD-GEN.
(1) Generation Phase: The POL simulation is executed using a predefined set of parameters specified in an external configuration file. Each simulation instance runs on a single CPU core; however, the framework supports multi-core parallelism by executing multiple independent simulation instances concurrently. The simulator also provides checkpointing functionality, allowing execution to pause at designated checkpoints and resume from saved states with modified parameters. This enables efficient exploration of alternative hypotheses by branching multiple parallel simulation worlds from a common baseline state.
(2) Calibration Phase: In this phase, key statistical metrics, implemented as scoring functions within the framework, are extracted from a reference real-world dataset and compared against the corresponding outputs of the simulation. A parallel genetic algorithm is implemented to optimize POL parameters by maximizing statistical similarity between simulated and observed data. Calibration is optional and may be skipped when fidelity to a reference dataset is not required or when users prefer to manually specify simulation parameters.
(3) Processing Phase: The simulator generates multiple log files in fixed-size 512 MB chunks, including agent state, check-in, and social link records. These files are concatenated, and only the required columns are extracted and transformed into structured datasets suitable for downstream analysis. During this process, the simulation coordinate reference system, for example EPSG:26782, is automatically converted to EPSG:4326 to ensure compatibility with standard GIS tools.
(4) Visualization Phase: In this phase, the processed datasets are used for mobility analysis and visual analytics. This includes extracting key patterns such as trips, stay-point insights, check-in attributes, and movement clusters, as well as visual exploration of individual agents. These capabilities enable users to better understand mobility behavior, evaluate model performance, and gain insights into dataset characteristics.
3.2. POL Details
This section describes the new features added to POL as part of this papers extension, and the changes made to the original POL simulation to improve its performance. Additionally, we explain the checkpointing, which allows for saving and resuming simulation states, and the log streaming feature that enables real-time monitoring of simulation progress.
3.3. Performance Optimization
Table 1. Performance comparison between Vanilla and Enhanced simulations for a 10 simulation day run. Initialization and simulation times are reported for different numbers of agents. A value of -1 indicates that the simulation did not complete. (Init: Initialization time; Sim: Simulation time; Exit On/Off: agent can/cannot exit; s: seconds; m: minutes; h: hours)
Agents Vanilla Exit On Vanilla Exit Off Enhanced Exit On Enhanced Exit Off
Init Sim Init Sim Init Sim Init Sim
1000 33.2 s 10.3 m 30.7 s 15.2 m 8.7 s 1.3 m 4.9 s 1.0 m
2000 1.9 m 36.2 m 1.4 m 51.5 m 11.6 s 2.0 m 11.1 s 2.0 m
3000 2.6 m 1.4 h 2.2 m 1.9 h 19.1 s 3.1 m 20.9 s 3.1 m
4000 4.2 m 3.0 h 4.1 m 3.6 h 35.3 s 4.8 m 35.6 s 5.3 m
5000 4.6 m 3.7 h 5.7 m 4.3 h 48.8 s 5.7 m 53.0 s 5.8 m
10000 13.9 m 15.8 h 13.0 m 16.7 h 3.3 m 14.7 m 3.3 m 14.7 m
15000 24.1 m 35.4 h 22.8 m 41.2 h 8.7 m 32.6 m 8.0 m 28.9 m
20000 -1 -1 -1 -1 12.8 m 40.7 m 11.2 m 36.9 m
25000 -1 -1 -1 -1 18.9 m 56.7 m 19.3 m 58.0 m
50000 -1 -1 -1 -1 1.1 h 2.7 h 1.0 h 2.7 h
100000 -1 -1 -1 -1 4.4 h 9.0 h 4.5 h 9.4 h
150000 -1 -1 -1 -1 9.6 h 19.3 h 9.8 h 20.1 h
As we aim to simulate larger populations within the Pattern of Life (POL) simulation, performance optimization becomes a critical concern. The original Vanilla POL simulation exhibited limitations in scalability, particularly as the number of agents increased. To address these challenges, we implemented a series of performance optimizations in the Enhanced POL simulation. The Enhanced simulation incorporates several key optimizations, including optimize the initial parameter values, avoid logging redundant or unnecessary information, and replace the optimal travel distance algorithm with a Euclidean distance calculation (Khan et al., 2024) to identify the nearest location. These changes collectively contribute to a more efficient simulation process, reducing both initialization and execution times.
To evaluate the effectiveness of the proposed performance optimizations, we focus on the execution time of a single simulation instance. Since parallel execution does not alter the runtime of an individual simulation, this evaluation isolates the impact of the internal optimizations introduced in the enhanced system. All experiments were conducted by simulating a period of 10 days (14,400 simulation ticks) while varying the number of agents. The experiments were executed on a machine equipped with an Intel Core i7-6700HQ CPU and 24 GB of RAM. We compare the baseline Vanilla simulation against the Enhanced simulation, which incorporates the proposed performance optimizations and we will use the Enhanced simulation as the default simulation in HD-GEN going forward unless otherwise specified.
For each configuration, we report two distinct performance phases:
  • Initialization time, which measures the duration required to load spatial data, instantiate agents, initialize environments, and prepare simulation state.
  • Simulation time, which measures the wall-clock time required to execute the full 10-day simulation period.
Results are reported for four execution modes: Vanilla and Enhanced simulations, each evaluated with agent exit conditions enabled (exit_on) and disabled (exit_off). The exit condition allows agents to leave the simulation when predefined criteria are met, reducing computational load in some scenarios. A value of -1 indicates that the simulation did not complete within a reasonable time frame or was manually terminated due to excessive runtime.
Table 1 summarizes the results across a wide range of agent counts. The results demonstrate that the Vanilla simulation fails to scale beyond approximately 15,000 agents, frequently exceeding practical execution limits. In contrast, the Enhanced simulation completes successfully for populations exceeding 150,000 agents, while substantially reducing both initialization and simulation times across all tested configurations. These results confirm that the proposed optimizations significantly improve scalability while preserving the behavioral semantics of the underlying simulation.
Check-pointing
Figure 4. Deterministic checkpointing and cross platform execution of the simulation, showing how a paused run can be resumed and branched across Windows, macOS, and Linux machines to evaluate multiple scenarios from a shared state.
The POL simulation is deterministic, meaning that executing the simulation with the same input parameters always produces identical results. The simulation is also portable across different machines and operating systems, allowing execution to continue without breaking the pipeline or altering outcomes. These properties are essential for exploring multiple scenarios while preserving previously executed configurations.
Many studies require evaluating alternative scenarios under identical baseline conditions. For example, in an infectious disease setting, one may vary the number of initially infected individuals or compare different vaccination policies while keeping all other parameters fixed. To support such analyses, the simulation provides a checkpointing mechanism, illustrated in Fig. 4. In this example, the simulation begins on a Windows machine and runs until a checkpoint is reached. At this point, execution is paused and the complete simulation state is serialized. From the same checkpoint, the simulation can be resumed with different configurations on multiple machines. For instance, one branch continues on a macOS machine with a modified configuration, while another resumes on Windows without reconfiguration. Additional branches are launched on Linux machines to evaluate further parameter variations. Each resumed execution represents a distinct scenario derived from the same initial state. All resulting runs generate logs that are stored in a unified format and remain fully compatible regardless of the execution platform. The underlying operating system is transparent to downstream processing, ensuring that all scenarios can be analyzed and compared consistently in subsequent analysis phases.
Log Streaming
Figure 5. Real time streaming architecture for POL simulation logs, showing how agent events are published during execution and consumed by external analytics frameworks through a read only streaming interface.
Another aspect of the simulated world is real time log streaming, which enables interaction with simulation outputs while execution is in progress, rather than waiting for the simulation to complete and writing logs to static files. The POL simulation acts as the engine that generates agent behavior and produces event logs as scenarios evolve. A streaming server built on top of the logging system provides real time access to simulation data and publishes agent level events such as agent identifiers, timestamps, locations, and states. As shown in Fig. 5, the simulator interacts with the POL model to satisfy agent needs and control behavior based on agent characteristics, while streaming events through a message broker to external systems. Downstream frameworks, including Hadoop, Spark, and PostgreSQL, can subscribe to these streams to consume simulation outputs in real time. The streaming interface follows a read only design and does not modify agent behavior or simulation state. Instead, it enables live observation, monitoring, and analysis of agent behavior as simulation time advances, supporting online analytics and early insight into emerging patterns.
Figure 6. Example script for generating a new simulation map using a bounding box and output path.
3.4. HD-GEN Details
Generation Phase (1)
To generate simulated data, we developed tools for regional map preparation, simulation execution, and large scale dataset generation in parallel. Map generation relies on OpenStreetMap (OSM) data accessed through the Overpass API (API, 2024) to extract building footprints, road networks, and land use information. These raw features are further refined using predictive models to classify building usage. The map generation process produces three shapefiles, buildings.shp, buildingUnits.shp, and walkways.shp, which collectively define the spatial environment used by the POL simulation. Figure 6 shows an example script that generates a new map for the Minneapolis region. In addition to bounding box based queries, the tools support custom Overpass queries, allowing flexible selection of geographic regions and features for map generation.
Figure 7. An example of a Python script to run multiple simulation instances in parallel
We also developed tools that allow users to define simulation parameters in a JSON file and run multiple instances in parallel, reusing calibration results when available. Two execution models are supported: batch-processing for dependent tasks (e.g., genetic algorithm tuning) and execution-line for independent runs (e.g., large-scale data generation). The number of parallel instances is configurable based on system capacity. In Fig. 7, the Python script demonstrates how to execute multiple simulation instances concurrently. The function get_configured_params loads parameters from initial.json, allowing users to customize more than 60 default settings. Simulations are launched using run() with parallel execution enabled (parallel=8).
When batch_processing=False, simulation instances are executed independently in an execution pipeline, which is suitable for large scale data generation. When batch_processing=True, simulations are executed in groups, and execution waits until all instances in a group complete, which is useful for calibration and parameter tuning. All results, including simulation outputs and execution metadata, are stored in params.simulated.json for subsequent analysis or reuse. This approach efficiently manages computational resources while preserving simulation integrity, enabling scalable data generation and systematic computational studies.
Figure 8. HD-GEN simulation generation phase architecture showing parallel configuration, task parallelism, and batch processing modes for scalable log generation.
Fig. 8 illustrates the generation phase architecture of HD-GEN. The process begins with an initial set of parameters, which are used to configure the simulation and create executable instances for each parameter set. Based on the value of the batch processing flag, simulation instances are dispatched to one of two execution modes. When batch processing is enabled, instances are sent to the batch processing unit, where simulations are executed in groups. Execution proceeds group by group, and each group waits until all simulations in that group complete before the next group is started. This mode is primarily used for calibration and coordinated experimentation. When batch processing is disabled, instances are sent to the task parallelism unit, where simulations are executed independently along an execution line. In this mode, execution is limited only by the parallel capacity of the system, making it suitable for large scale data generation. In both modes, all simulation runs produce logs that are collected and stored as the final output of the generation pipeline.
Calibration Phase (2)
Figure 9. Genetic algorithm workflow for obtaining optimized parameters. The initial parameter pool evolves through mutation (Producer) and mixing strategies (Mixer) to form new generations. After simulation and score evaluation, weak parameter sets are removed while strong ones are retained. The process continues until convergence, and the highest scoring parameters are selected
For calibration, we used the GeoLife (Zheng et al., 2010, 2011) dataset as a real-world reference and applied a genetic algorithm guided by a similarity function to tune the simulation parameters. To keep the example simple, we compared real and simulated data using four metrics: average distance per trip (ADT), average distance per agent per day (ADA), maximum trip distance (MXD), and median trip distance (MDD). The similarity score in Equation 1 quantifies the match, with values closer to 1 indicating greater similarity to the GeoLife dataset.
(1) $\text{Similarity}(G,P)=1-\frac{1}{|M|}\sum_{k\in M}\frac{|k(P)-k(G)|}{k(G)}$
Here, $G$ is the GeoLife dataset, $P$ is the simulated dataset, M=ADT,ADA,MXD,MDDM={ADT,ADA,MXD,MDD}, and $k(\cdot)$ denotes the value of the metric.
To calibrate the simulations, we developed a genetic algorithm integrated into the data generation framework. We identified 63 parameters affecting agent behavior (e.g., number of interests, walking speed, rental cost ratio). A full list and documentation are provided in our GitHub repository ({https://github.com/onspatial/hd-gen}). The goal was to find parameter values producing trajectories that best match GeoLife metrics. The algorithm begins with simulations of $n$ using valid parameter values randomly chosen. Each run is scored using Equation 1, and the best performing parameter sets are retained as “parents.” New parameter sets (“children”) are then generated by combining parent values in five ways: maximum, minimum, mean, random mix, or random mutation. This process repeats until $n$ children are produced, each followed by a simulation and scoring step. New generations are iteratively created until convergence or manual termination. The best performing parameters across all generations are selected as the calibration result. The implementation is publicly available on GitHub.
Fig. 9 illustrates the genetic algorithm process used to obtain optimized parameters. We start with an initial set of parameters that are placed into a pool. Two methods are then applied to generate new parameters:
  1. (1)
    Producer: Creates new parameters by applying mutations. It uses the valid range defined for each parameter. For example, if the parameter joviality must be between 0 and 1, the producer will only generate values within that range.
  2. (2)
    Mixer: Combines existing parameters using different strategies:
    • Max: selects the maximum value among the inputs. For example, if the values for a parameter $x$ are $1,2,3,$ and $4$, the mixed value becomes $4$.
    • Min, Mean, and Random apply their respective rules in a similar manner.
The results from the producer and mixer form the new generation. From this generation, a fixed number of parameter sets (equal to the layer size) are selected for the next stage. After running simulations, each parameter set receives a score using the Equation 1. Any set with a score below a defined threshold is removed, and the surviving parameter sets are returned to the pool. The algorithm continues iterating until a stopping condition is reached. Finally, the highest scoring parameter sets, such as the top one or top ten, are chosen as the optimized parameters.
Processing Phase (3)
Figure 10. Post simulation log processing workflow (processing phase), illustrating column selection, data integration, conversion, and downstream usage for analysis, labeling, or calibration.
The raw outputs generated by the simulator need to be transformed into structured, application ready datasets which is done through a modular and scalable processing pipeline. After data generation, simulation logs are processed in two primary ways: (1) by concatenating the complete set of split log files, or (2) by trimming and merging logs from a selected subset of simulation instances. In both cases, standardization is often required. For example, spatial coordinates recorded in the agentStateTable are not stored in standard GPS format and must be converted before downstream use. Our processing scripts convert the raw logs into structured datasets suitable for tasks such as machine learning model training, calibration, and statistical analysis. The agentStateTable, which records the state of each agent at every simulation tick, can be split into manageable chunks, filtered by relevant fields, and reassembled into a unified dataset. Users can explicitly specify which fields are retained, enabling targeted data extraction tailored to specific research objectives.
As illustrated in Fig. 10, post simulation processing begins with column selection, followed by integration of log fragments into a consolidated dataset. A data conversion step then transforms fields into formats required by downstream applications. The resulting dataset can be directly analyzed and visualized, labeled for supervised learning tasks, or passed to the scorer during the calibration phase without further modification.
HD-GEN is designed with scalability and customizability as core principles. To reduce unnecessary log generation, we provide detailed documentation on our Github page that guides users in configuring the simulator to emit only essential fields such as timestamps, geographic coordinates, and agent identifiers. This design minimizes storage and computational overhead while improving overall efficiency. For very large datasets, the pipeline includes Python scripts capable of processing hundreds of gigabytes of data, as well as lightweight Bash utilities that perform file system level concatenation and column extraction. In addition to filtering redundant fields, the pipeline supports embedding ground truth labels and other annotations, ensuring that generated datasets are well aligned with downstream research and application requirements.
Visualization Phase (4)
Another critical aspect of HD-GEN is the visualization of processed data. Visualization offers a powerful means of interpreting simulation outputs by uncovering patterns, anomalies, and behavioral dynamics that are often difficult to detect in raw numerical logs (Grimm, 2002). It is particularly valuable for validating the effects of parameter adjustments, identifying unexpected or emergent agent behaviors, and improving the overall accuracy of the model. When new features or anomalous behaviors arise, visual inspection provides an intuitive method for confirming their presence and assessing their impact across both temporal and spatial dimensions.
To facilitate these analyses, we provide a comprehensive visualization toolkit specifically designed for simulation logs. This toolkit supports a wide range of visualization modes, including time-series plots, spatial trajectory maps, and aggregated statistical summaries. By enabling researchers to rapidly explore simulation results, the toolkit enhances both debugging efficiency and interpretability of complex agent interactions. Together, these visualization capabilities complement the data processing pipeline, ensuring that raw outputs can be transformed not only into structured datasets but also into interpretable visual representations that accelerate validation and insight generation.
4. Generated Dataset
In this section, we present example datasets that were generated and processed using our software and have been previously published. Datasets smaller than 100 GB are hosted on OSF.io, with access links provided through our GitHub repository. For datasets exceeding 100 GB, we instead provide detailed instructions in the GitHub documentation to guide researchers in reproducing the data locally. This approach ensures that even very large datasets remain accessible while also allowing users to scale the simulation further by increasing the number of agents or extending the simulation duration. Most datasets described in this work were generated using a compact desktop machine equipped with a 2.40 GHz Intel i5-1135G7 processor (8 cores, 16 GB RAM) running Linux Fedora. This demonstrates that the framework is efficient and does not require high-performance computing resources to produce large-scale mobility datasets, making it broadly accessible to both academic and industry researchers.
4.1. Geolife+: Calibrated Dataset
To demonstrate the scalability of our framework, we generated multiple datasets based on the GeoLife trajectories, varying both the number of simulated agents and the simulation duration. These datasets range from small-scale runs with hundreds of agents to large-scale scenarios involving tens of thousands of agents, spanning months to several years. Compared to the original GeoLife dataset, our simulations capture a richer set of agent activities, as every agent is continuously observed throughout the simulation period. This produces datasets that are significantly larger in size while maintaining realistic mobility characteristics, offering researchers an expanded view of population-scale mobility (Amiri et al., 2024c).
4.2. Massive Trajectory Data
We generated a collection of large-scale check-in, location-based social network (LBSN) and trajectory datasets across four diverse urban and suburban regions: Fairfax (George Mason University campus area), New Orleans (French Quarter), San Francisco, and Atlanta. Each dataset integrates three components: simulated check-ins, evolving social links, and trajectory points recorded at five-minute intervals for every agent. Simulations were conducted at varying scales, from 1,000 to 10,000 agents over 15-month periods, as well as long-term runs spanning up to 20 years with 1,000 agents. These datasets capture both short-term dynamics and long-term behavioral patterns, providing a flexible foundation for mobility, social network, and anomaly detection research (Amiri et al., 2023).
4.3. Human Mobility Dataset with Anomalies
Figure 11. Example of visualization phase output for simulated check-in data with global anomalies introduced over a 1050-day period
To demonstrate the flexibility of HD-GEN, we generated datasets that incorporate controlled anomalies designed specifically for anomaly detection research. The simulation environment allows users to inject a wide range of behavioral deviations into agent activities, including irregular eating patterns, atypical social interactions, and disruptions to routine work behavior. Each anomaly can be configured at different intensity levels and applied either to individual agents or to the entire population, enabling the creation of diverse and realistic experimental scenarios. Agents can be designated as anomalous through three complementary mechanisms: (i) random assignment, in which anomalies are stochastically applied to selected agents; (ii) infectious disease style propagation, where anomalous behavior spreads dynamically through agent interactions, mimicking epidemiological processes; and (iii) location based triggers, where agents begin exhibiting anomalies after visiting specific locations. These mechanisms support the study of both isolated and systemic anomaly patterns and enable investigation into how abnormal behaviors emerge, propagate, and are detected in complex mobility systems (Amiri et al., 2024b).
In related work, we introduced specific anomaly types to generate datasets for global anomaly detection tasks (Zhang et al., 2024a, b; Liu et al., 2024). The implemented anomaly categories include hunger anomalies, social anomalies, and work anomalies. Hunger anomalies cause agents to experience hunger more frequently than normal, resulting in increased check ins at restaurants or homes. Social anomalies cause agents to ignore usual preferences and social influences, selecting recreational locations at random. Work anomalies disrupt routine behavior by causing agents to skip work on days when attendance is expected. Each anomaly type is further parameterized by three intensity levels: red, orange, and yellow. Red level anomalies represent extreme behavioral deviations, such as skipping work 100% of the time, orange level anomalies represent moderate deviation at approximately 50%, and yellow level anomalies represent mild deviation at approximately 20%. Anomalies can be applied either to specific agents for a defined duration or globally across all agents on selected days.
Using the visualization tools described in 3.4, Fig. 11 presents an example of a global anomaly experiment. The figure shows check in patterns for multiple venue types, including apartments, pubs, restaurants, and workplaces, over a 1050 day simulation period starting in January 2024. The first 450 days simulate normal behavior, allowing agents to establish stable baseline routines. During the remaining 600 days, global anomalies are introduced, where all agents adopt a specific anomaly type and intensity level on each day. As illustrated in the figure, agent behavior remains stable during the baseline phase and becomes increasingly irregular once anomalies are introduced. For example, during red level global hunger anomaly events, all agents experience heightened hunger, leading to increased visits to restaurants and homes. This produces sharp spikes in restaurant and home check ins, accompanied by fluctuations in workplace visits as agents leave work to eat.
We used the provided tools to generate the dataset reported in (Amiri et al., 2024b), which contains multiple anomaly types and configurations. The framework further supports flexible control over anomaly distribution by allowing users to select between random assignment, infectious propagation, and location based spreading mechanisms. Together, these capabilities enable the generation of rich, realistic datasets for evaluating anomaly detection methods across a wide range of behavioral and mobility scenarios.
5. Conclusions
In this paper, we have presented a software system to generate, calibrate and process large-scale synthetic geospatial datasets by simulating the pattern of life of individuals with the consideration of real-world constraints. We have shown the effectiveness of our approach by generating a set of synthetic datasets based on the GeoLife dataset. Furthermore, we have shown that the generated datasets exhibit similar statistical properties to the original dataset, and can be used for various geospatial data analysis tasks. We have also provided detailed instructions on how to reproduce the generated datasets, and have made the code and data available on GitHub. With the provided datasets and code, researchers can easily generate large-scale synthetic geospatial datasets for their research purposes and evaluate the performance of their algorithms on realistic data. Our approach bridges the gap between real-world and simulated trajectory data by integrating statistical features from real datasets with the scalability of simulations. By introducing a data generation tool, a genetic algorithm for calibration, and a data processing pipeline, we enable the creation of realistic, large-scale mobility datasets suitable for diverse applications. This framework provides a robust foundation for research in human mobility modeling, machine learning, and data-driven applications while addressing the limitations of both real and synthetic datasets.
6. Acknowledgments
Funding
Supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0025. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.
This research was also supported by the National Science Foundation under Award Abstract #2109647, Data-Driven Modeling to Improve Understanding of Human Behavior, Mobility, and Disease Spread.
Use of Language Assistance Tools
We clarify that to write this paper we used ChatGPT 5.1 only to improve the English and readability of our writing. We did not ask it to write any sections or paragraphs from its own knowledge.